Braeden's father is 4 times as old as Braeden is. In 20 years his father will be only twice as old as Braeden. How old are Braden and his father now
Answer by krutarthas(22)

Q:-Braeden's father is 4 times as old as Braeden is. In 20 years his father will be only twice as old as Braeden. How old are Braden and his father now ?
ans:- let us take Braeden age is x year
Braeden's father is 4 times as old as Braeden is
so Braeden's father age is 4x year
after 20 year Braeden age is x+20 year
after 20 year Braeden father age is 4x+20 year
but given condition is --In 20 years his father will be only twice as old as Braeden's age means
4x+20=2(x+20)
⇒4x+20=2x+40
⇒4x-2x=40-20
⇒2x=20
⇒x=20/2=10
Braeden age is x year= 10 years
Braeden's father age is 4x year= 4X10 = 40 years
answer
Krutartha